OpenStudy (anonymous):

Add following polynomial: Show all steps PLEASE.(1/5x^2 - 1/5x - 3/4) + (1/10x^2 - 1/2x + 1/2)

OpenStudy (anonymous):

I'll get you started: You can only add (combine) "like terms". In this case, the two x^2 terms add together, the two x terms add together, and the two constants add together. Now you have to remember the rules for adding fractions by getting a common denominator. So, to add 1/5x^2 and 1/10x^2, you want to need a common denominator for 1/5 and 1/2, the lowest one being 10. 1/5x^2+1/10x^2 = 2/10x^2 + 1/10x^2 = 3/10x^2 Repeat the process to add -1/5x+(-1/2x) and -3/4+1/2. Other notes: Always simplify results if you can, and if the problem was subtraction instead of addition, you must remember to use the distributive property.